Item4 1points Return to question Item 4 Item 4 1 points Treasury bills are currently paying...

Item4

1points

Return to question

Item 4

Item 4 1 points

Treasury bills are currently paying 8 percent and the inflation rate is 3 percent.

a.

What is the approximate real rate of interest? (Enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)

b. What is the exact real rate? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
Answer #1

a)Approximate real rate of interest = Nominal rate of interest - Inflation rate

                                      = 8 -3

                                      = 5%

b)Exact real rate =[(1+ nominal rate)/(1+inflation rate) ] - 1

             = [(1+.08)/ (1+.03) ] -1

            = [1.08 /1.03 ] -1

            = 1.0485 - 1

            = .0485 or 4.85%
Answer #2

ANSWER 


a.


Approximate real rate

= Treasury rate - Inflation rate

= 8% - 3%

= 5% (ANSWER).


b.


Exact real rate 

= (1 +Treasury rate) / (1 + inflation rate) - 1

= (1 + 0.08) / (1 + 0.03) - 1

= 0.0485

= 4.85%  (ANSWER).

answered by: Tulsiram Garg